# Summarizing Long Regulatory Documents with a Multi-Step Pipeline

### v1

@@ 2 Related work
The idea of multi-step methods is to leverage both extractive and abstractive techniques to alleviate the burden of summarising very long documents. Pilault et al. Pilault et al. (2020) add one extractive step before generating the abstractive summary. The extractive parts are then used beside the original text as input for the transformer. A related approach is taken in CreativeSumm Kim et al. (2022) for the summarisaiton of length movie scripts. Liu et al. (2018) summarise Wikipedia articles by first performing an extractive step, using the extracted sentences as additional input to the summariser. Bleiweiss (2023) propose a two-step method for long biographical novels. Klaus et al. (2022) make use of a two-step method to summarize legal regulatory documents. Klaus et al. use TextRank Mihalcea and Tarau (2004), a graph-based extractive summarization approach, for the first extractive step and BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019) for a second extractive step.
@@ 3.3 Evaluation
Due to resource and time constraints, we selected specific architectures to be included in the qualitative evaluation. To analyse the impact of different extractive models, we compare different versions of BART, using (1) the best extractive model; (2) no extractive step; (3) the best legal LM for extraction; and (4) the best long-context extractive model To analyse the impact of different abstractive strategies, we also include (5) the best long-context abstractive model; and (6) the best decoder-only model.
@@ 4 Results
Extractive model Ratio type ROUGE-1 ROUGE-2 ROUGE-L BERTScore BARTScore BLANC
@@
Abstractive model Ratio type ROUGE-1 ROUGE-2 ROUGE-L BERTScore BARTScore BLANC
@@
Extr. model Ratio type Abstr. model Fact. correctness Usability Accuracy Fluency Coherence
1 RoBERTa Dependent BART 2.0 2.0 1.5 1.5 2.0
2 - No extraction BART 3.5 1.0 2.0 3.0 1.5
3 LexLM Dependent BART 4.0 3.5 3.0 3.0 3.0
4 Longformer Dependent BART 2.0 2.0 2.5 1.5 2.0
5 - No extraction PegasusX 3.5 1.0 2.5 3.0 1.0
6 RoBERTa Dependent Llama3 3.0 2.5 2.5 2.5 2.0
Table 5: Average human evaluation results.
@@ 4.2 Comparison of abstractive models
For every abstractive model, two versions are compared: one leveraging RoBERTa with a dependent ratio and one without using any extractive step at all. The results for all abstractive models and their variants can be seen in Table 4. For clarity, models that incorporate an extractive step will be referred to by the name of the abstractive model. Models that do not use an extractive step will be denoted by appending “-NE” to the name of the abstractive model, where “NE” signifies “No Extraction.”
@@ 4.3 Human evaluation
There are interesting divergences between the expert evaluation and the conclusions drawn based the automatic metrics, an observation which is quite common in the NLG and summarisation literature (cf. Belz and Reiter, 2006; Reiter, 2018; Celikyilmaz et al., 2021). In particular, experts suggest that legal LMs help achieve more satisfactory summaries if used in the extractive step. On the other hand, both automatic and human evaluation suggest that BART is a competitive model for summarisation, especially if preceded by an extractive step.
@@ 5 Conclusion
In this paper, we focused on summarisation of long regulatory documents. Our findings indicate that while models with a longer context length do not benefit from extraction, an extractive step renders BART, an encoder-decoder architecture, highly competitive. Human experts confirm this finding, but also indicate a preference for summaries relying on extraction with a domain-specific, legal language model.
Future work should consider whether these findings are generalisable to other domains. Furthermore, a more extensive human evaluation is required to ensure that our findings are reliable.
@@ Appendix B Human evaluation results
Architecture # Extr. model Ratio Abstr. model Fact. correctness Usability Accuracy Fluency Coherence
1 RoBERTa Dependent BART 1 1 1 1 1
2 - No extraction BART 3 1 1 2 1
3 LexLM Dependent BART 4 3 3 2 2
4 Longformer Dependent BART 1 1 1 1 1
5 - No extraction PegasusX 4 1 2 1 1
6 RoBERTa Dependent Llama3 3 1 2 1 1
Table 8: Human evaluation results participant 1.
Architecture # Extr. model Ratio Abstr. model Fact. correctness Usability Accuracy Fluency Coherence
1 RoBERTa Dependent BART 3 3 2 2 3
2 - No extraction BART 4 1 3 4 2
3 LexLM Dependent BART 4 4 3 4 4
4 Longformer Dependent BART 3 3 4 2 3
5 - No extraction PegasusX 3 1 3 5 1
6 RoBERTa Dependent Llama3 3 4 3 4 3
Table 9: Human evaluation results participant 2.

### v2

@@ 2 Related work
The idea of multi-step methods is to leverage both extractive and abstractive techniques to alleviate the burden of summarising very long documents. Pilault et al. Pilault et al. (2020) add one extractive step before generating the abstractive summary. The extractive parts are then used beside the original text as input for the transformer. A related approach is taken in CreativeSumm Kim et al. (2022) for the summarisation of lengthy movie scripts. Liu et al. (2018) summarise Wikipedia articles by first performing an extractive step, using the extracted sentences as additional input to the summariser. Bleiweiss (2023) propose a two-step method for long biographical novels. Klaus et al. (2022) make use of a two-step method to summarize legal regulatory documents. Klaus et al. use TextRank Mihalcea and Tarau (2004), a graph-based extractive summarization approach, for the first extractive step and BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019) for a second extractive step.
@@ 3.3 Evaluation
Due to resource and time constraints, we selected specific architectures to be included in the qualitative evaluation. To analyse the impact of different extractive models, we compare different versions of BART, using (1) the best extractive model; (2) no extractive step; (3) the best legal LM for extraction; and (4) the best long-context extractive model. To analyse the impact of different abstractive strategies, we also include (5) the best long-context abstractive model; and (6) the best decoder-only model.
@@ 4 Results
Extractive model Ratio type R1 R2 RL BERTScore BARTScore BLANC
@@
Abstractive model Ratio type R1 R2 RL BERTScore BARTScore BLANC
@@
Extr. model Ratio Abstr. model FC U Acc Fl Coh
1 RoBERTa Dep. BART 2.0 2.0 1.5 1.5 2.0
2 - NE BART 3.5 1.0 2.0 3.0 1.5
3 LexLM Dep. BART 4.0 3.5 3.0 3.0 3.0
4 Longformer Dep. BART 2.0 2.0 2.5 1.5 2.0
5 - NE PegasusX 3.5 1.0 2.5 3.0 1.0
6 RoBERTa Dep. Llama3 3.0 2.5 2.5 2.5 2.0
Table 5: Average human evaluation results. Dep: Dependent ratio; NE: No extraction; FC: Factual Correctness; U: Usability; Acc: Accuracy; Fl: Fluency; Coh: Coherence
@@ 4.2 Comparison of abstractive models
For every abstractive model, two versions are compared: one leveraging RoBERTa with a dependent ratio and one without using any extractive step at all. The results for all abstractive models and their variants can be seen in Table 4. For clarity, models that incorporate an extractive step will be referred to by the name of the abstractive model. Models that do not use an extractive step will be denoted by appending “-NE” to the name of the abstractive model, where “NE” signifies “No Extraction”.
@@ 4.3 Human evaluation
A somewhat surprising outcome is that LLaMA-3 scores relatively poorly on coherence and fluency, compared to the best-performing model. It should be noted that the two evaluators diverged significantly in their scores for this model on these criteria (compare Tables 8 and 9 in Appendix B). Furthermore, as noted above, LLaMA was treated somewhat differently since it is the only decoder-only model. In particular, we subtracted 1500 tokens from the model’s context length to accommodate the extractive summary; this too could have impacted results, though we adjusted the extractive summarisation process to ensure minimal truncation.
Despite the fact that this is a small-scale evaluation (a point we return to in Section 5), there are interesting divergences between expert judgments and the conclusions drawn based on the automatic metrics, an observation which is quite common in the NLG and summarisation literature (cf. Belz and Reiter, 2006; Reiter, 2018; Celikyilmaz et al., 2021).
In particular, experts suggest that legal LMs help achieve more satisfactory summaries if used in the extractive step. On the other hand, both automatic and human evaluation suggest that BART is a competitive model for summarisation, especially if preceded by an extractive step.
@@ 5 Conclusion
In this paper, we focused on summarisation of long regulatory documents. Our findings indicate that while models with a longer context length do not benefit from extraction, an extractive step renders BART, an encoder-decoder architecture, highly competitive. A small-scaled evaluation with human experts confirms this finding. However, experts also indicate a preference for summaries relying on extraction with a domain-specific, legal language model.
Future work should consider whether these findings are generalisable to other domains. Furthermore, a more extensive human evaluation is required to ensure that our findings are reliable. This is particularly crucial given that human expert judgments are not perfectly aligned with the outcomes of our metric-based evaluation, which echoes findings from other studies. A further possible research direction is to use a state-of-the-art LLM as an evaluator or ‘judge’ for generated texts, a strategy which recent research suggests is increasingly viable Liu et al. (2023); Zheng et al. (2023), though also one that requires some caution in view of results suggesting self-bias on the part of LLMs Panickssery et al. (2024), as well as lower reliability in comparison with expert judgment Bavaresco et al. (2024).
@@ Appendix B Human evaluation results
Architecture # Extr. model Ratio Abstr. model FC U Acc Fl Coh
1 RoBERTa Dep. BART 1 1 1 1 1
2 - NE BART 3 1 1 2 1
3 LexLM Dep. BART 4 3 3 2 2
4 Longformer Dep. BART 1 1 1 1 1
5 - NE PegasusX 4 1 2 1 1
6 RoBERTa Dep. Llama3 3 1 2 1 1
Table 8: Human evaluation results participant 1. Dep: Dependent ration; NE: No extraction; FC: Factual Correctness; U: Usability; Acc: Accuracy; Fl: Fluency; Coh: Coherence
Architecture # Extr. model Ratio Abstr. model FC U Acc Fl Coh
1 RoBERTa Dep. BART 3 3 2 2 3
2 - NE BART 4 1 3 4 2
3 LexLM Dep. BART 4 4 3 4 4
4 Longformer Dep. BART 3 3 4 2 3
5 - NE PegasusX 3 1 3 5 1
6 RoBERTa Dep. Llama3 3 4 3 4 3
Table 9: Human evaluation results participant 2. Dep: Dependent ration; NE: No extraction; FC: Factual Correctness; U: Usability; Acc: Accuracy; Fl: Fluency; Coh: Coherence